OpenStudy (anonymous):

A spherical storage tank has a diameter of 14 ft. How many cubic feet of water will it hold? (Use = .)

OpenStudy (diyadiya):

use volume of sphere =\[4/3 \pi r ^{3}\]

OpenStudy (amistre64):

im guessing it will use the volume of a sphere formula

OpenStudy (amistre64):

\[\frac{4pi\ r^3}{3}\]

OpenStudy (amistre64):

\[2pi\int_{0}^{14}(14^2-x^2)dx\]right?

OpenStudy (diyadiya):

\[(4/3)\pi 7^{3}\]= 1436.0266...

OpenStudy (amistre64):

doh!! .... forgot to adjust for diameter to radius ...